# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

### v3

@@ Abstract
By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost. However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real-world settings.
Our empirical study reveals that some experts encode redundant knowledge during pre-training. We thus propose a method of grouping and pruning similar experts to improve the model’s parameter efficiency. We validate the effectiveness of our method by pruning three state-of-the-art MoE architectures, including Mixtral, Deepseek-MoE, and Qwen. The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. We will release our code to facilitate future research.
@@
Zeliang Zhang1111This research was conducted during the author’s internship at Microsoft Research.  Xiaodong Liu2   Hao Cheng2  Chenliang Xu1  Jianfeng Gao2 1University of Rochester 2Microsoft Research {zeliang.zhang, chenliang.xu}@rochester.edu, {xiaodl, cheng.hao, jfgao}@microsoft.com
@@ 1 Introduction
There have been several studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. (2022) prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. (2024) find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. (2024a) merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. (2024) explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. (2024) enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
@@
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 3, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for three state-of-the-art MoE architectures, including Mixtral (Mixtral-8x7B and Mixtral-8x22B), Deepseek-MoE (Deepseek-MoE-16B) and Qwen (Qwen2-57B-14A). The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. Our contribution can be summarized as follows,
@@ 2 Related Work
3 Methodology
Notations. We denote $F(\cdot;\Theta,W,K)$ as an MoE layer, where $\Theta=\{\theta_{1},\theta_{2},...,\theta_{N}\}$ indicates the parameter set for $N$ experts $\{f_{n}(\cdot;\theta_{n})\}_{n=1}^{N}$, $W\in\mathbb{R}^{N\times d}$ indicates the routing weights for each expert and $K$ indicates the number of experts selected for each token.
For a new token $x\in\mathbb{R}^{d}$, we first compute the routing logits to score the match between $x$ and the $n$-th expert:
$p_{n}(x)=\frac{e^{W_{n}x}}{\sum_{t=1}^{N}e^{W_{t}x}},$ (1)
where $W_{t}\in\mathbb{R}^{d}$ is the routing policy of $t$-th expert. Then, we take the experts $\{f_{n}(\cdot;\theta_{n})\}_{n=i_{1}}^{i_{K}}$ with top-$K$ logits $\{l_{n}\}_{n=i_{1}}^{i_{K}}$ as the selected target experts. Last, we compute the weighted average of the outputs from selected experts as the final output of the MoE layer,
$y=\sum\limits_{n=i_{1}}^{i_{K}}p_{n}(x)\cdot f_{n}(x;\theta_{n}).$ (2)
@@ 3.1 Task Definition
Previous works (Lu et al., 2024; He et al., 2024; Li et al., 2024a) have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
Specifically, for the $l$-th MoE layer $F^{l}(\cdot;\Theta^{l},W^{l},K)$ in the given LLM $\mathcal{M}(\cdot;\mathcal{F})$, we want to prune $r$ experts from the total of $N$, $l=1,2,...,L$. This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete. We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ towards the following optimization objective:
$\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y),$ (3)
where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the new LLM with pruned MoE layer $\hat{\mathcal{F}}$, and $\mathcal{D}$ is any dataset under the task-agnostic setting.
This task is particularly challenging, as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions. In this work, we propose to estimate the expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
@@
3.2 Expert Similarity in MoE layer
Motivation: Intuitively, the more independent the experts are within the same MoE layer, the more diverse the representations learned by each expert. This diversity makes it harder to replace or drop any of the experts. Conversely, the redundancy of shared knowledge learned by different experts provides opportunities for model pruning.
We employ the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) as the criteria to evaluate the similarity between experts in each MoE layer. Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
$\rho_{ij}=\frac{\text{HSIC}(K^{i},K^{j})}{\text{HSIC}(K^{i},K^{i})\cdot\text{HSIC}(K^{j},K^{j})},$ (4)
where $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function.
Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 (Raffel et al., 2020). We respectively use the linear kernel and RBF kernel in similarity computation of eq. 4, where the use of linear kernel actually corresponds to the dot product similarity between the expert outputs.
As illustrated in Fig. 2, the darker the cell, the more similar the two related experts. The two kernel-based distance measures demonstrate similar changes in expert similarity across different layers. There are varying moderate ratios of similar paired experts in different MoE layers, suggesting the plausibility of pruning the MoE layer with many experts down to fewer experts. For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance. Full evaluation results can be found in appendix B.
3.3 Discover and Merge Similar Experts on Feature Space
@@
Having identified the existence of moderate expert similarity in each MoE layer, we propose pruning the experts within each MoE layer by merging experts with shared knowledge. We present an overview of our method in fig. 3. Suppose we want to reduce the number of experts from $N$ to $r$, the details of our approach are outlined as follows.
Discovering Similar Experts from the Expert Graph. We construct a unidirectional graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ to group similar experts into the same clusters, where we denote $\mathcal{V}$ as the set of nodes and $\mathcal{E}$ as the connections between nodes.
First, for each expert $f_{i}$, we compute its feature representation $\mathcal{M}(f_{i})$, where $\mathcal{M}(\cdot)$ is a specialized function to represent $f_{i}$.
Second, we compute the pairwise distance based on expert representation and obtain the disparity matrix $D\in\mathbb{R}^{N\times N}$, where $D_{ij}=\text{DistanceFunction}(\mathcal{M}(f_{i}),\mathcal{M}(f_{j}))$.
Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the pair-wise expert distance matrix $D$.
@@
$\min\sum_{i=1}^{r}\left(\sum_{j,k\in\mathcal{V}_{i}}D_{jk}-\sum_{t\neq i}^{r}\sum_{j\in\mathcal{V}_{i},k\in\mathcal{V}_{t}}D_{jk}\right),$ (5)
@@
where The objective function seeks to minimize the intra-group similarity (first term) and maximize the inter-group difference (second term) based on the pairwise similarity $D_{jk}$ based on expert $i$ and $j$.
Merging Similar Experts with Routers. To diversity the experts and preserve the different knowledge learned by different models clustered in the same group (Wortsman et al., 2022; Wang et al., 2022), we merge the clusted experts with their routers on the weight space as follows,
@@
where $\mathcal{V}_{n}$ is the set of similar experts in the $n$-th cluster $\mathcal{G}_{n}$, and we have $\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}=1$. We update the MoE layer $F$ by respectively replacing all the experts grouped in $\mathcal{G}_{n}$ by a single FFN expert layer $f(\cdot;\hat{\theta}_{n})$ and corresponding routing weights $\hat{W}_{n}$, $n=1,2,...,N-r$.
3.4 Practical Implementation
On the discovery step. We propose two strategies for computing the disparity matrix $D$, including the data-centric strategy and model-centric strategy.
For the data-centric strategy, since the full pre-training datasets for different models are large and inaccessible, we use the C4 dataset, a commonly used smaller pre-training subset that serves as an effective surrogate for capturing task-agnostic knowledge of experts Lu et al. (2024). Specifically, by disabling the router function, for the same input, we use the output of each expert as its expert representation $\mathcal{M}(\cdot)$. To enhance generalization and mitigate overfitting to the selected samples during model pruning, we apply data augmentation by randomly mixing token embeddings during the representation computation at the discovery stage.
For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation. One is to leverage the vectorized weights directly ($\clubsuit$). The other is to leverage the local linearity of neural networks (Zhang and Wu, 2020) to compute the surrogate weight matrix ($\spadesuit$).
@@ 3.4 Practical Implementation
On the merging step. We consider three approaches in practice deciding $\alpha$ to merge similar grouped experts. The first is to only maintain the experts with the maximum visiting frequency and drop all the others. The second is to set $\alpha_{i}=\frac{1}{|\mathcal{V}_{n}|}$, which uniformly assembles all the experts grouped. The last one is learning $\alpha$ to merge the grouped experts by minimizing the following loss function,

### v4

@@ Abstract
In this work, we address the memory overhead of deploying Mixture-of-Experts (MoE) architectures in Large Language Models (LLMs). While MoE layers improve LLM performance without increasing inference costs, the ever-growing number of experts inflates memory requirements, hindering practical deployment. Our empirical study reveals that some experts encode redundant knowledge during pre-training. We thus propose a method of grouping and pruning similar experts to improve the model’s parameter efficiency. We validate the effectiveness of our method by pruning three state-of-the-art MoE architectures, including Mixtral, Deepseek-MoE, and Qwen. The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
@@
thanks: Work done during the internship at Microsoft Research.Zeliang Zhang1  Xiaodong Liu2   Hao Cheng2  Chenliang Xu1  Jianfeng Gao2 1University of Rochester 2Microsoft Research {zeliang.zhang, chenliang.xu}@rochester.edu, {xiaodl, cheng.hao, jfgao}@microsoft.com
@@ 1 Introduction
There have been several studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. (2022) prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. (2024) find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. (2024a) merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. While effective, these methods depend on knowing the target tasks. In contrast, task-agnostic pruning methods that do not rely on task information are more appealing and useful in real-world applications because they can apply to both seen and unseen tasks. However, this is more challenging since there are no explicit task and data cues to guide which experts are redundant. He et al. (2024) explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. (2024) enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
@@
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 3, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for three state-of-the-art MoE architectures, including Mixtral (Mixtral-8x7B and Mixtral-8x22B), Deepseek-MoE (Deepseek-MoE-16B) and Qwen (Qwen2-57B-14A). The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. Our contributions are summarized as follows,
@@ 2 Related Work
3 Method
In this section, we present our approach for pruning experts within MoE layers in large language models. Our goal is to identify experts that share highly similar knowledge and merge them, thereby reducing the model size and improving efficiency without significantly degrading performance.
3.1 Notation and Preliminaries
We denote $F(\cdot;\Theta,W,K)$ as an MoE layer. Here, $\Theta=\{\theta_{1},\theta_{2},...,\theta_{N}\}$ is the set of parameters for $N$ experts $\{f_{n}(\cdot;\theta_{n})\}_{n=1}^{N}$, where each expert is typically a feed-forward network (FFN). The routing matrix $W\in\mathbb{R}^{N\times d}$ determines which experts are selected for each token. K denotes how many top experts are chosen for each input token.
For a token $x\in\mathbb{R}^{d}$, we first compute the routing logits to measure how well it matches each expert: $p_{n}(x)=\frac{e^{W_{n}x}}{\sum_{t=1}^{N}e^{W_{t}x}}$. Then, we select top-$K$ experts, $\{f_{n}(\cdot;\theta_{n})\}_{n=i_{1}}^{i_{K}}$, based on these logits. Finally, we compute the MoE layer’s output as the weighted combination of the chosen experts: $y=\sum\limits_{n=i_{1}}^{i_{K}}p_{n}(x)\cdot f_{n}(x;\theta_{n}).$
@@ 3.2 Task Definition
Previous works (Lu et al., 2024; He et al., 2024; Li et al., 2024a) have shown that pruning some experts in MoE layers can improve model efficiency in both inference and fine-tuning without causing model collapse. However, an open question remains: which experts should be removed?
Formally, consider a LLM $\mathcal{M}(\cdot;\mathcal{F})$ comprsing $L$ MoE layers. Each MoE Layer $F^{l}(\cdot;\Theta^{l},W^{l},K)$ has $N$ experts. Our objetive is to only remain $r$ experts from each MoE layer:
$\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{N-r}}\}$ (1)
$\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{N-r}}\}$
We aim to find these indices $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{N-r}\}$ that minimizes loss on a generic dataset $D$:
$\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y).$ (2)
The search space $(C^{r}_{N})^{L}$ is enormous, making direct combinatorial search intractable.
3.3 Measuring Expert Similarity with CKA
@@
Our key insight is that experts exhibiting similar behaviors likely contain redundant knowledge. By identifying and pruning these redundant experts, we can simplify model and improve its efficiency.
To quantify similarity, we use Centered Kernel Alignment (CKA) (Kornblith et al., 2019; Davari et al., ; Smerkous et al., ) as the criteria to evaluate the similarity between experts in each MoE layer. Intuitively, CKA measures how similarly two experts transform a shared batch of inputs. For any two experts $f_{i}$ and $f_{j}$, given a batch of inputs $x=\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ is computed as follows,
$\rho_{ij}=\text{CKA}(K^{i},K^{j})=\frac{\text{HSIC}(K^{i},K^{j})}{\text{HSIC}(K^{i},K^{i})\cdot\text{HSIC}(K^{j},K^{j})},$ (3)
where $K^{i}$ and $K^{j}$ are kernel matrices constructed from experts’ outputs on the same input batch $x$, and HSIC is the Hilbert-Schmidt Independence Criterion (Greenfeld and Shalit, 2020). The definition is as follows: $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function. Notably, all experts are provided with the same input batch—no token distribution by the router function is involved—allowing for a clearer representation of the distinct knowledge each expert acquires within the same layer.
We evaluate the similarity between experts in Mixtral-8Bx7B using 32 randomly selected samples from the C4 pre-training dataset (Raffel et al., 2020). Following eq. 3, we compute expert similarity with both a linear kernel (equivalent to using a dot product) and an RBF kernel.
As shown in fig. 2, darker cells indicate greater similarity between pairs of experts. Both kernel measures reveal similar patterns across layers, indicating that some experts are moderately to highly similar. This suggests the feasibility of pruning certain MoE layers. For example, in the first layer, experts 2 through 5 consistently show similarity scores above 0.7, suggesting that removing one of these experts would likely have minimal impact on overall model performance. Full evaluation results can be found in appendix B.
3.4 Discovering and Merging Similar Experts
As shown in fig. 3, our pruning strategy consists of two key steps to reduce the number of experts from $N$ to $r$: (1) identifying groups of similar experts and (2) merging each group into a single expert.
@@ 3.4 Discovering and Merging Similar Experts
Discovering Similar Experts from the Expert Graph. We construct a unidirectional graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{A}\}$ to cluster similar experts, where $\mathcal{V}$ is the set of nodes (each representing an expert), $\mathcal{E}$ denotes the set of edges connecting experts that exhibit positive similarity, and $\mathcal{A}$ is a weight matrix that encodes these pairwise similarities. The procedures for constructing $\mathcal{A}$ is shown as follows.
First, we represent each expert $f_{i}$ by its output embedding on a shared calibration dataset (data-centric) or by its own weights (model-centric) if data is unavailable, as $\mathcal{R}(f_{i})$.
Second, we calculate the weight on each edge of $\mathcal{E}$ to represent the similarity between experts, i.e., $\mathcal{A}\left(\mathcal{E}\right)$, using CKA (data-centric) as eq. 3 or mean squared error (model-centric)***More discussion is provided in Appendix G. Each element $\mathcal{A}\left(\mathcal{E}_{ij}\right)=\text{CKA}(\mathcal{R}(f_{i}),\mathcal{R}(f_{j}))$ reflects how similar experts $f_{i}$ and $f_{j}$ are in the transformed space $\mathcal{R}(f)$ as the previous step.
@@
$\max\sum_{i=1}^{r}\left(\sum_{j,k\in\mathcal{V}_{i}}\mathcal{A}\left(\mathcal{E}_{jk}\right)-\sum_{t\neq i}^{r}\sum_{j\in\mathcal{V}_{i},k\in\mathcal{V}_{t}}\mathcal{A}\left(\mathcal{E}_{jk}\right)\right),$ (4)
@@
Our objective is to minimize the intra-group similarity (first term) and maximize the inter-group difference (second term) based on the pairwise similarity $\mathcal{A}\left(\mathcal{E}_{ij}\right)$ based on expert $i$ and $j$.
Merging Similar Experts To diversify the experts and preserve the different knowledge learned by different models clustered in the same group (Wortsman et al., 2022; Wang et al., 2022; Zhao et al., 2024), we merge the clustered experts with their routers on the weight space as follows,
@@
where $\mathcal{V}_{n}$ is the set of similar experts in the $n$-th cluster $\mathcal{G}_{n}$, and we have $\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}=1$. We update the MoE layer $F$ by respectively replacing all the experts grouped in $\mathcal{G}_{n}$ by a single FFN expert layer $f(\cdot;\hat{\theta}_{n})$ and corresponding routing weights $\hat{W}_{n}$, $n=1,2,...,N-r$. The pseudo-code is provided in appendix G.
3.5 Practical Considerations
Computing the disparity matrix $\mathcal{A}(\mathcal{E})$. We propose two strategies for computing the disparity matrix $\mathcal{A}(\mathcal{E})$: a data-centric strategy and a model-centric strategy.
For the data-centric strategy, since the full pre-training datasets for different models are large and inaccessible, we use the C4 dataset, a commonly used smaller pre-training subset that serves as an effective surrogate for capturing task-agnostic knowledge of experts Lu et al. (2024). Specifically, by disabling the router function, for the same input, we use the output of each expert as its expert representation $\mathcal{R}(\cdot)$. Then, we compute the expert similarity and discover the similar experts by eq. 4. To enhance generalization and mitigate overfitting to the selected samples during model pruning, we apply data augmentation by randomly mixing token embeddings during the representation computation at the discovery stage.
For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation. One is to leverage the vectorized weights directly ($\clubsuit$). The other is to leverage the local linearityIn short, the ’local linearity’ property of neural networks refers to the observation that, despite having multiple hidden layers and non-linear activation functions, neural networks exhibit linear behavior within a local region. This property motivates our design of a single-layer approximation using the surrogate weight. of neural networks (Zhang and Wu, 2020) to compute the surrogate weight matrix ($\spadesuit$).
@@ 3.5 Practical Considerations
The selection of distance function. While the CKA metric provides an accurate estimation of expert similarity, it can be both memory- and computation-intensive. As an alternative, cosine similarity may be used as an efficient approximation. This corresponds to computing CKA for a single sample (data-centric), replacing the $K_{i}$ and $K_{j}$ in eq. 3 with the weights of the $i$-th and $j$-th experts, respectively (model-centric). In practical deployments, we also observe that mean squared error performs well across several models, including DeepSeek-MoE and Qwen-MoE.
Discussion on two strategies. Ensuring distributional similarity between surrogate and real datasets is essential for the data-centric method to accurately estimate expert similarity; without the guarantee, its performance could be degraded significantly. While pre-trained data for many public MoEs, such as Mixtral, DeepSeek, and Qwen, is unavailable, our study, along with prior research, demonstrates that leveraging a small subset of the open-source dataset C4 can still achieve competitive performance. Alternatively, the model-centric approach presents a robust solution, eliminating the need for data access during pruning and, in some cases, achieving performance comparable to or exceeding that of data-centric methods. More discussion is provided in appendix H.
Merging Strategies and Trade-offs. We consider three approaches in practice deciding $\alpha$ to merge similar grouped experts. The first is to only maintain the experts with the maximum visiting frequency and drop all the others. The second is to set $\alpha_{i}=\frac{1}{|\mathcal{V}_{n}|}$, which uniformly assembles all the experts grouped. The last one is learning $\alpha$ to merge the grouped experts by minimizing the following loss function,